# A simple explanation of the absence of the spherical nuclei with the hidden-color states

### v3

@@ Abstract
Abstract: Inspired by the EMC effect, the Cd puzzle and the SU3-IBM, a hypothesis can be given for a nucleus, that only the nucleus itself is a trivial (0,0) representation of the SU(3) group, which leads to the simple conclusion that spherical nucleus does not exist and the spherical mean field is not allowed. The key conclusion is that the color degree of freedom is required even at the low energy excitation of the nucleus, and the whole nucleus should have many hidden-color states.
@@
The past decade has seen some perspective shifts in the field of nuclear physics. An important one is the discovery of the Cd puzzle Wood11 , which makes the traditional collective mode of phonon excitation of the nucleus with spherical shape to be questioned and refuted Heyde16 ; Garrett18 . On the basis of these related experiments Garrett08 ; Garrett10 ; Garrett12 ; Batchelder12 ; Garrett19 ; Garrett20 , I proposed that this is a new mode of collective excitation, and found that the SU3-IBM can be used to explain this new mode Wang22 ; WangPt ; Wang24 . The three papers lead us to question the traditional perspective on the emergence of the collective behaviors in the nucleus. Some global effects in the nucleus may be neglected in previous studies.
The prominent SU(3) symmetry appears in the SU3-IBM, for which one will doubt whether this will be related to the SU(3) symmetry in the quantum chromodynamics Drischler21 ; Gross23 ; Achenbach24 . In existing studies, especially after the discovery of the EMC effect Arneodo94 ; Thoma95 ; Weise00 ; Hen17 , some researchers began to understand some phenomena in the nucleus from the quark-gluon degree of freedom Bakker14 ; Sargsian03 ; Clot19 . But no one believed that the low-energy excitation of the nucleus is related to the color degree of freedom. In this Letter, I establish such a connection for the first time. The results of these experiments for the absence of the spherical nuclei can have a very simple explanation. While some researchers have discussed the hidden-color state in the nucleus Harvey81 ; Harvey812 ; Ji86 ; Abbas86 ; Myhrer88 ; Ping11 ; Miller14 ; Entem20 , no one thought that the whole nucleus itself has many hidden-color states.
@@
No one believed that there may be other possibility until the Cd puzzle experiments denied this idea Garrett08 ; Garrett10 ; Garrett12 ; Batchelder12 ; Garrett19 ; Garrett20 . The coupling between the normal states and the intruder states is in fact weak Garrett12 , and the spherical phonon excitation of the Cd nuclei is refuted Heyde16 ; Garrett18 ; Wood11 . The normal states present a new spherical-like $\gamma$-soft collective mode Wang22 ; WangPt ; Wang24 . This means that the origin of the collective nature in the nucleus requires an alternative explanation. Just as Heyde and Wood said: “The emerging picture of nuclear shapes is that quadrupole deformation is fundamental to achieving a unified view of nuclear structure.” and “a shift in perspective is needed: sphericity is a special case of deformation.” Heyde16
The Cd puzzle challenges the conventional view of nuclear structure, because these experimental results are more than expected. Inspired by the Cd puzzle, I proposed the extended interacting boson model Iachello87 ; Casten10 ; Wang08 with SU(3) higher-order interactions (SU3-IBM) Wang22 , just satisfying the idea of Heyde and Wood. In this model, only the U(5) limit and the SU(3) limit are included. It suggests that the SU(3) symmetry dominates the onset of the quadrupole deformation wang23 ; zhou23 . Thus not only the quadrupole deformation but also the SU(3) symmetry is important for the unified description of nuclear structure. The Cd puzzle really presents a new spherical-like deformation Wang22 ; WangPt ; Wang24 .
@@
Therefore, the current situation is that both experiment and theory are beginning to support the conclusion that the spherical nucleus does not exist. But it also raises new problems. There are two important conclusions in the SU3-IBM: (1) higher-order interactions are important and the third-order interaction can be even dominant Wang22 ; wang23 ; Wang , (2) the SU(3) symmetry governs all the interactions except for the pairing force (the $d$ boson number operator). Thus the idea of the mean field requires some corrections. The two-body interaction may not necessarily dominate the residual ones. And an important question is, where does this SU(3) symmetry in the SU3-IBM come from?
@@
Obviously, the success of the SU3-IBM proves that the key to the Cd puzzle, the fact that the spherical nucleus does not exist, is related to the SU(3) symmetry. This result is very interesting. In the SU(3) representation, (0,0) is the spherical shape. Thus the Cd puzzle implies that the shape of the nucleus can not be the (0,0) representation of the SU(3) group. In this Letter, a simple but strict explanation on this impossibility is given.
The Cd puzzle is the low-energy phenomenon of the nucleus, which is related to the collective or global effect of the nucleons in the nucleus. It is hard to imagine that this phenomenon may be related to the high-energy behavior. However, in the nucleus, especially for the low-energy excitation, the only one with the SU(3) symmetry is the color symmetry of the quantum chromodynamics Drischler21 ; Gross23 ; Achenbach24 , so the two must be linked together. Here some unexpected relations are found. The EMC effect Arneodo94 ; Thoma95 ; Weise00 ; Hen17 means that quarks and gluons still play an important role in the nucleus. It is found that the color degrees of freedom is also required in the low-energy excitation of the nucleus.
@@
The EMC effect illustrates that the result of deep inelastic scattering is related to the target of the used nucleus, and the effect increases as the mass number of the nucleus increases Arneodo94 ; Thoma95 ; Weise00 ; Hen17 . This means that the quark-gluon distribution of the nucleons in the nucleus is not the same as the ones in the free nucleons, which is related to the whole nuclei. It thus provides an example that the large-scale nucleus can affect the quark-gluon distribution in the small-scale nucleons. Recently, researchers have generally agreed that this phenomenon is related to the medium modification of the nucleus, fundamentally, to the color confinement Hen17 .
@@
Therefore in the nucleus, except for the nucleons and various instantaneous mesons, quarks and gluons need to be considered. There are two possibilities for the existence of this quark-gluon degree of freedom in nuclei. One is that although the nucleons in the nucleus change, the quarks and gluons are still bound in the nucleons. Another is, in the nucleus, quarks and gluons are not only bound within the nucleons, but also occasionally appear between the nucleons. For the latter, the nucleons in the nucleus may not be color-neutral, and they have color. Of course, due to the color confinement, these colored nucleons must form a color-neutral whole entity, which is called the hidden-color state. Thus some researchers guess that “multi-quark hidden-color components exist in ordinary nucleus” Bakker14 ; Close82 ; Gabathuler85 .
@@
Researchers have also studied the effect of the quarks and gluons in the nucleus Sargsian03 ; Clot19 , and the hidden-color state Bakker14 ; Harvey81 ; Harvey812 ; Ji86 ; Abbas86 ; Myhrer88 ; Ping11 ; Miller14 ; Entem20 ; Miller19 ; Miller20 . Just as Cloët et al. said:“ it is unlikely that the nucleon-meson based approaches can remain valid or contain the correct degrees of freedom for all processes at all energy scales. Clearly identifying these scales and processes is key to exposing the role of quarks and gluons in nuclei and thereby developing an understanding of how nuclei emerge within QCD.” Clot19 The quark-gluon degree of freedom is needed in the discussion of nuclei. The key is to find clear or revelatory evidence. It’s hard to imagine that quarks and gluons are only confined to the interior of the nucleons and mesons while the nucleus has a larger volume. In this Letter I find the relationship between the absence of a spherical nucleus and the hidden-color states of the entire nucleus.
@@
Quarks and gluons have color degrees of freedom which has SU(3) symmetry. The color confinement is a conjecture that has not been proved by theory. It means that composite particles made up of quarks and gluons, such as hadrons or nuclei, are color-neutral, or the (0,0) representations of SU(3) group. Here I focus on the nuclei. Hydrogen nucleus is the proton, so it can be expressed by the (0,0) representation of the SU(3) group. The same is true for the neutron, and also true when extended to nuclei with larger mass number $A$. The whole composite of the quarks and gluons must be color-neutral. The colour confinement is a global phenomenon.
@@
This is a bold step forward understanding the color degree of freedom in the nucleus in this Letter. Although this is to be expected, there is a lack of evidence to suggest this in previous studies.
@@
This assumption certainly needs the theoretical confirmation (color confinement), but here we will see that it directly gives the conclusion that a spherical nucleus does not exist. It has now been confirmed that the nuclear force can be divided into three parts: the mean field, short-range strong correlation, and long-range correlation Hen17 . The mean field presents the single particle effect, which means that the nucleons in the nucleus have some of the same features. The short-range strong correlation presents the correlation between only the two nucleons, and only the long-range correlation contains the collective or global effect among the whole nucleons in the nucleus.
Recent experiments have also demonstrated that the EMC effects are associated with short-range strong correlations, but the more comprehensive correlations with the three parts remains unclear. It is already known that the EMC effects can be related to the corrections of the mean-field nucleons and short-range strongly correlated nucleons, but the exact mechanism is still unclear Hen17 ; Hen11 ; Hen12 . Thus some researchers will doubt whether such medium correction will be related with the low-energy nuclear structure. Recently Machleidt and Sammarruca said:“ it is a good question to ask whether medium modifications of nuclear forces show up in a noticeable way and/or are even needed for quantitative nuclear structure predictions.” and asked: “Are the energies typically involved in conventional nuclear structure physics low enough to treat nucleons as structure-less objects?” Here I give the answer, even at the low-lying excitation, it is not possible. The Cd puzzle means the nucleon modifications and the color degree freedom is also needed if the assumption is correct.
@@
In most previous studies, the color degree of freedom is not considered, that is, the nucleons in the nucleus are color-neutral, then $(\lambda_{k},\mu_{k})$ in equation (1) are all (0,0). However if this equation (1) is a hidden-color state, it can not be (0,0). This is the focus of this Letter. It should be stressed here that the intermediate coupling result can not appear as the (0,0) representation. When $A$ increases many coupling possibilities can be obtained for the hidden-color state, so there are in fact many hidden-color states with the mass number $A$ for the whole nucleus. All these possibilities in equation (1) will be divided into the mean field, short-range correlations and long-range correlations. Therefore, these three parts are related to each other, and they cannot be expressed by (0,0) due to the existence of the hidden-color states. Although the way of the division is still unclear, the result is clear.
@@
Detailed studies should also consider other degrees of freedom of the nucleons, such as orbital, isospin and spin Bakker14 . Since this is not a specific calculation, there is no careful consideration. This Letter only considers about the color degree of freedom for the color confinement.
@@
One point to be emphasized here is that the color correlation with more nucleons must be weaker, and why such a hidden-color state associated with the entire nucleus appears. The possibility is that the more nucleons this state is related to, the bigger the number of such states will be. This requires further theoretical confirmation. In Ref. Bakker14 , it is shown that, the number of the hidden-color states with 9 quarks is much more than the one with 6 quarks Bakker14 . In a nucleus, the hidden-color states with various nucleon numbers may exist, but the most important may be the two-body correlation (the strongest), and the overall correlation caused by the hidden-color state of the whole nucleus (the most possibilities). This is exactly what the experiment has been found out to be.
@@
Since there are many hidden-color states in the nucleus, a possible result will occur. These hidden-color states will have quantum interference with each other, and if destructive interference occurs, the long-range correlation may disappear and the magic number appears. If constructive interference occurs, a large deformation will occur. This is what to be focusing on in the future. If the magic numbers can be explained based on this result, then it will further prove the correctness of this conclusion here.
One might argue that the state of the nucleon in the nucleus is not definite and is a possible superposition state, such as
$|\psi\rangle=\alpha_{0}|(0,0)\rangle+\alpha_{1}|(\lambda_{1},\mu_{1})\rangle+\cdot\cdot\cdot+\alpha_{k}|(\lambda_{k},\mu_{k})\rangle+\cdot\cdot\cdot,$ (2)
@@
here $\alpha_{k}$ is the coefficients of the kth superposition state. This is of course possible. However, due to the color confinement, the key to such a state is still to generate equation (1). Just as Bakker and Ji said:“ Nuclear systems are identified as color-singlet composites of quark and gluon fields, beginning with the six-quark Fock component of the deuteron. An immediate consequence is that nuclear states are a mixture of several color representations which cannot be described solely in terms of the conventional nucleon, meson, and isobar degrees of freedom: there must also exist hidden-color multi-quark wave-function components, i.e., nuclear states which are not separable at large distances into the usual color-singlet nucleon clusters.” The nucleons or their combinations may have the color-neutron component (0,0), but they can not the (0,0) representation. In this Letter, I give the first evidence that this hidden-color state should exist.
@@
Second the long-range correlations are not (0,0). This long-range correlations lead to the deformations of the nucleus. If they are not (0,0), it means that it cannot be spherical. This is exactly what the Cd puzzle is trying to reveal. This conclusion is also supported when the Cd puzzle is verified by the SU3-IBM. In the new spherical-like $\gamma$-soft mode, the SU(3) third-order Casimir operator is as important as the second-order one. This is a new kind of deformation. In Ref. wang23 , the shape phase transition from the prolate shape to the oblate shape in the Hf-Hg region is studied, and the ratio of the SU(3) second-order and third-order Casimir operators is shown in Fig. 1. This value decreases approximately linearly as the number of nucleons increases (the number of bosons decreases). Recently it is also found that the boson number odd-even effect Fortunato11 ; Zhang12 of the 196-204Hg really exists, in which the SU(3) third-order interaction plays the main role Wang . This means that when the valence nucleons increase, the two-body residual interaction may not dominate. This is not quite the usual mean field viewpoint. I guess that the multiple possibilities of the color confinement in the nucleus in equation (1) let the low-energy two-body interactions interfere destructively, reducing it. The appearance of the magic numbers may arise from the complete quantum interference between the hidden-color states. For the Cd puzzle can be well explained by the SU3-IBM, the hidden-color state in equation (1) is much more important than expected, and the (0,0) component in equation (2) may be nearly ignored for the collective phenomenon.
@@
This hypothesis is strongly supported by the EMC effect, which needs to be further verified. The Cd puzzle is the direct consequence of this hypothesis. Although this explanation is easy, it fits with the results of the experiments. The conclusion that spherical nucleus does not exist can be confirmed by the Cd puzzle experiments and can be supported by other surrounding nuclei, such as Te, Pd nuclei Garrett18 . The results of these experiments can be well described by the SU3-IBM Wang22 ; WangPt ; Wang , which further supports this hypothesis.
@@
The result is somewhat unexpected. For the first time, it explicitly points out a clear correlation between the low and high energy behaviors of the nucleus, and while the true reason for the correlation is still unknown, it does exist. Of course, it can be guessed that this correlation comes from the quark-gluon effect in the nucleus, which would otherwise not explain the SU(3) symmetry in the SU3-IBM. These results illustrate that the color degree of freedom is influential even in the field of low-energy nuclear structure. This can be supported by the SU(3) symmetry of the SU3-IBM.
So this hypothesis is supported by both experiment and theory. The role of the quarks and gluons in the nucleus, though not clear, exists and works at low energy excitation. This has been overlooked in previous studies. Of course, this is inevitable. Because the change of the quarks and gluons in the nucleus is the behaviors found in the high-energy region, it is difficult to expect that the EMC effect can affect the low-energy nuclear structure. This requires support from the experiments. The Cd puzzle, as well as its SU3-IBM description, becomes the possible evidence.
@@
In conclusion, a simple explanation for the absence of the spherical nuclei is given based on the EMC effect, the short-range strong correlations, the Cd puzzle and other new experimental results, the calculation of the SU3-IBM and the hidden-color state. It can be obtained that the Cd puzzle is the result of the color confinement in the nucleus. If the hidden-color states exist, the spherical nucleus can not appear. It is clearly pointed out that the color degree of freedom is necessary even in low energy excitation of nuclear structure. The correlations between the low-energy and high-energy behaviors in the nucleus does exist, and this is the first time to be found here, but what it is requires more further studies.
In this Letter a new theoretical framework to understand nuclear forces is established phenomenologically and a possible big picture for the whole nuclear physics is drawn. Since this explanation is very simple and rigorous, the new connection deserves serious consideration. It does not conflict with previous effective field theories. The possibility of the color degree of freedom between the nucleons is considered here. The SU3-IBM supports this possibility and gives a guide for future theoretical developments. In future, the possible results of this new perspective will be discussed.

### v4

@@ Abstract
Abstract: Inspired by the EMC effect, the Cd puzzle and the SU3-IBM, a hypothesis can be given for a nucleus, that only the nucleus itself is a trivial (0,0) representation of the SU(3) group, which leads to the simple conclusion that spherical nucleus does not exist and the spherical mean field is not allowed. The key conclusion is that the whole nucleus should have many hidden-color states, and the quark-gluon degrees of freedom are required even at the low-energy excitation of the nucleus.
@@
The past decade has seen some perspective shifts in the field of nuclear physics. An important one is the discovery of the Cd puzzle Wood11 , which makes the traditional perspective on the collective mode of phonon excitation of the nucleus with spherical shape to be questioned and refuted Heyde16 ; Garrett18 . On the basis of these related experiments Garrett08 ; Garrett10 ; Garrett12 ; Batchelder12 ; Garrett19 ; Garrett20 , I proposed that this is a new mode of collective excitation, and found that the extended interacting boson model Iachello87 ; Casten10 ; Wang08 with SU(3) higher-order interactions (SU3-IBM) can be used to explain this new mode Wang22 ; WangPt ; Wang24 . These findings lead us to question the traditional perspective on the emergence of the collective behaviors in the nucleus. Since the Cd puzzle conflicts with previous ideas and the SU3-IBM is very different from previous nuclear structure theories, it can be inferred that some global effects in the nucleus may be neglected in previous studies and the emergence of the collectivity in the nucleus may have an implicit origin.
The prominent SU(3) symmetry appearing in the SU3-IBM provides a unique clue, for which one will doubt whether this will be related to the color SU(3) symmetry in the quantum chromodynamics Drischler21 ; Gross23 ; Achenbach24 . In existing studies, especially after the discovery of the EMC effect Arneodo94 ; Thoma95 ; Weise00 ; Hen17 , some researchers began to understand some phenomena in the nucleus from the quark-gluon degree of freedom, and especially the nuclear forces at short distances may require the quark-gluon interactions to explain their repulsion Bakker14 ; Abbas17 ; Sargsian03 ; Sargsian14 ; Clot19 . But no one believed that the low-energy excitation of the nucleus is related to the color degree of freedom. In this Letter, I establish such a connection for the first time. If the hidden-color states exist Harvey81 ; Harvey812 ; Ji86 ; Abbas86 ; Myhrer88 ; Ping11 ; Miller14 ; Miller19 ; Miller20 ; Entem20 , the results of these experiments for the absence of the spherical nuclei can have a very simple explanation. While some researchers have discussed the hidden-color state in the nucleus with fewer nucleons, no one was sure that the whole nucleus itself should have many hidden-color states. This implies that the shape deformation of the nucleus may arise from the quark-gluon short-distance interactions between the nucleons in the nucleus. Although this conclusion is somewhat different from previous ideas, it is actually reasonable.
@@
No one believed that there may be other possibility until the Cd puzzle experiments denied this idea Garrett08 ; Garrett10 ; Garrett12 ; Batchelder12 ; Garrett19 ; Garrett20 . The coupling between the normal states and the intruder states is in fact weak Garrett12 , and the spherical phonon excitation of the Cd nuclei is refuted Heyde16 ; Garrett18 ; Wood11 . Just as Garret, Green and Wood said: “it is possible that the Cd isotopes may not represent vibrational systems and that the essential physics of their motion has been missing” Garrett08 . I found that the normal states of the Cd nuclei present a new spherical-like $\gamma$-soft collective mode Wang22 ; WangPt ; Wang24 . This means that the origin of the collective nature in the nucleus requires an alternative explanation. Just as Heyde and Wood said: “The emerging picture of nuclear shapes is that quadrupole deformation is fundamental to achieving a unified view of nuclear structure.” and “a shift in perspective is needed: sphericity is a special case of deformation.” Heyde16
The Cd puzzle challenges the conventional view of nuclear structure, because these experimental results are more than expected. Inspired by the Cd puzzle, I proposed the SU3-IBM Wang22 , just satisfying the idea of Heyde and Wood. In this model, only the U(5) limit and the SU(3) limit are included. It suggests that the SU(3) symmetry dominates the onset of the quadrupole deformation wang23 ; zhou23 . Thus not only the quadrupole deformation but also the SU(3) symmetry is important for the unified description of nuclear structure. The Cd puzzle really presents a new spherical-like deformation Wang22 ; WangPt ; Wang24 .
@@
Therefore, the current situation is that both experiment and theory are beginning to support the conclusion that the spherical nucleus does not exist. But it also raises new problems. There are two important conclusions in the SU3-IBM: (1) higher-order interactions are important and the third-order interaction can be even dominant Wang22 ; wang23 ; Wang , (2) the SU(3) symmetry governs all the interactions except for the pairing force (the $d$ boson number operator). Thus the idea of the mean field requires some corrections that the two-body interaction may not necessarily dominate the residual ones. And an important question is, where does this SU(3) symmetry in the SU3-IBM come from?
@@
Obviously, the success of the SU3-IBM proves that the key to the Cd puzzle, the fact that the spherical nucleus does not exist, is related to the SU(3) symmetry. This result is very important for a further deeper understanding of the origin of the Cd puzzle. In the SU(3) representation, (0,0) presents the spherical shape. Thus the Cd puzzle implies that the shape of the nucleus can not be the (0,0) representation of the SU(3) group. In this Letter, a simple but strict explanation on this impossibility is given.
The Cd puzzle is the low-energy phenomenon of the nucleus, which is related to the collective or global effect of the nucleons in the nucleus. It is hard to imagine that this phenomenon may be related to the high-energy behavior. However, in the nucleus, especially for the low-energy excitation, the only one with the exact SU(3) symmetry is the color symmetry in the quantum chromodynamics Drischler21 ; Gross23 ; Achenbach24 , so the two must be linked together. Although somewhat beyond expectations, as the SU3-IBM was proposed to resolve the Cd puzzle, this is the unique possibility. Here some unexpected relations are found. The EMC effect Arneodo94 ; Thoma95 ; Weise00 ; Hen17 means that quarks and gluons still play an important role in the nucleus. It is found that the color degrees of freedom is also required in the low-energy excitation of the nucleus.
@@
The EMC effect illustrates that the result of deep inelastic scattering is related to the target of the used nucleus, and the effect increases as the mass number of the nucleus increases Arneodo94 ; Thoma95 ; Weise00 ; Hen17 . This means that the quark-gluon distribution of the nucleons in the nucleus is not the same as the ones in the free nucleons, which is related to the whole nucleus. It thus provides an example that the large-scale nucleus can affect the quark-gluon distribution in the small-scale nucleons. Just as Miller said: “At least three length scales (the radius of the nucleus, the average inter-nucleon separation distance, and the size of the nucleon) are important in gaining a complete understanding of the physics of nuclei” and “the three scales are closely related, so that a narrow focus on any given specific range of scales may prevent an understanding of the fundamental origins of nuclear properties” Miller20 . Recently, researchers have generally agreed that this phenomenon is related to the medium modification of the nucleus, fundamentally, to the color confinement Hen17 .
@@
Therefore in the nucleus, except for the nucleons and various instantaneous mesons, quarks and gluons need to be considered. There are two possibilities for the existence of this quark-gluon degree of freedom in nuclei. One is that although the nucleons in the nucleus change, the quarks and gluons are still bound in the nucleons. Another is, in the nucleus, quarks and gluons are not only bound within the nucleons, but also occasionally appear between the nucleons. For the latter, the nucleons in the nucleus may not be color-neutral, and have color. Of course, due to the color confinement, these colored nucleons must form a color-neutral whole entity. Since these quark-gluon interactions can only occur within the short distances between the nucleons, the united state with many colored nucleons in the nucleus is called the hidden-color state Abbas17 . Thus some researchers guess that “multi-quark hidden-color components exist in ordinary nucleus” Bakker14 ; Close82 ; Gabathuler85 .
@@
Researchers have also studied the effect of the quarks and gluons in the nucleus Bakker14 ; Abbas17 ; Sargsian03 ; Sargsian14 ; Clot19 , and the hidden-color state Bakker14 ; Harvey81 ; Harvey812 ; Ji86 ; Abbas86 ; Myhrer88 ; Ping11 ; Miller14 ; Entem20 ; Miller19 ; Miller20 . Just as Cloët et al. said:“ it is unlikely that the nucleon-meson based approaches can remain valid or contain the correct degrees of freedom for all processes at all energy scales. Clearly identifying these scales and processes is key to exposing the role of quarks and gluons in nuclei and thereby developing an understanding of how nuclei emerge within QCD.” Clot19 The quark-gluon degree of freedom is needed in the discussion of nuclei. The key is to find clear or revelatory evidence. It’s hard to imagine that quarks and gluons are only confined to the interior of the nucleons and mesons while the nucleus has a larger volume. In this Letter I find the relationship between the absence of a spherical nucleus and the hidden-color states of the entire nucleus. For the existence of the Cd puzzle, the multi-quark hidden-color states should exist in ordinary nucleus.
Since the quark-gluon interactions can be only found within the short distances between the nucleons to produce a strong repulsive force that prevents the nucleons in the nucleus from collapsing into a much smaller sphere, that is, to maintain the stability of the nucleus Bakker14 ; Abbas17 ; Sargsian14 , this color force as the cause of the shape of the nucleus can be understandable. Without this effect, the shape of the nucleus would be meaningless.
@@
Quarks and gluons have color degrees of freedom which has SU(3) symmetry. The color confinement is a conjecture that has not been absolutely proved by theory. It means that composite particles made up of quarks and gluons, such as hadrons or nuclei, are color-neutral, or the (0,0) representations of the SU(3) group. Here I focus on the nuclei. Hydrogen nucleus is the proton, so it can be expressed by the (0,0) representation of the SU(3) group. The same is true for the neutron, and also true when extended to nuclei with larger mass number $A$. The whole composite of the quarks and gluons must be color-neutral. The colour confinement should be a global phenomenon Abbas17 , which is the reason for the existence of the hidden-color state.
@@
The quark-gluon interactions are very necessary when discussing the short distance repulsions between the nucleons Bakker14 ; Abbas17 ; Sargsian03 ; Sargsian14 ; Clot19 . The hidden-color states has been also discussed in Bakker14 ; Harvey81 ; Harvey812 ; Ji86 ; Abbas86 ; Myhrer88 ; Ping11 ; Miller14 ; Entem20 ; Miller19 ; Miller20 . This conclusion is obvious.
@@
This assumption certainly needs the theoretical confirmation (color confinement), but here we will see that it directly gives the conclusion that a spherical nucleus does not exist. It has now been confirmed that the nuclear force can be divided into three parts: the mean field, short-range strong correlation, and long-range correlation Hen17 . The mean field presents the single particle effect, which means that the nucleons in the nucleus have some of the same features. The short-range strong correlation presents the correlation between only the two nucleons, and only the long-range correlation contains the collective or global effect among all the nucleons in the nucleus.
Recent experiments have also demonstrated that the EMC effects are associated with the short-range strong correlations, but the more comprehensive correlations with the three parts remains unclear. It is already known that the EMC effects can be related to the corrections of the mean-field nucleons and short-range strongly correlated nucleons, but the exact mechanism is still unclear Hen17 ; Hen11 ; Hen12 . Thus some researchers will doubt whether such medium correction will be related with the low-energy nuclear structure. Recently Machleidt and Sammarruca said:“ it is a good question to ask whether medium modifications of nuclear forces show up in a noticeable way and/or are even needed for quantitative nuclear structure predictions.” and asked: “Are the energies typically involved in conventional nuclear structure physics low enough to treat nucleons as structure-less objects?” Here I give the answer, even at the low-lying excitation, it is not possible. The Cd puzzle means the nucleon modifications and the color degree freedom is also needed if the assumption is correct.
@@
In most previous studies, the color degree of freedom is not considered, that is, the nucleons in the nucleus are color-neutral, then $(\lambda_{k},\mu_{k})$ in equation (1) are all (0,0). However if this equation (1) is a hidden-color state, they can not all be (0,0). Here the case that each $(\lambda_{k},\mu_{k})$ is not (0,0) is considered, which is the focus of this Letter. It should be stressed here that the intermediate coupling result can not appear as the (0,0) representation too. When $A$ increases many coupling possibilities can be obtained for the hidden-color state, so there are in fact many hidden-color states with the mass number $A$ for the whole nucleus. All these possibilities in equation (1) will be divided into the mean field, short-range correlations and long-range correlations. Therefore, these three parts are related to each other, and they cannot be expressed by (0,0) due to the existence of the hidden-color states. In particular, the case that neither $(\lambda_{k},\mu_{k})$ is (0,0) and the intermediate coupling is also not (0,0) is the most common situation Bakker14 ; Abbas17 . Although the way of the division is still unclear, the result is clear. A simple case for all nuclei with $A\geq 2$ is that the nucleons in the hidden-color state are represented as (1,1), and their couplings are also (1,1), which makes this conclusion true.
@@
Detailed studies should also consider other degrees of freedom of the nucleons, such as orbital, isospin and spin Bakker14 . Since this is not a specific calculation, there is no careful consideration. This Letter only considers about the color degree of freedom for the color confinement. In this Letter, I do not present a specific relationship between the hidden-color states and the SU(3) representation of the shapes of the nuclei, because this is very complex and may even be very difficult at the moment. I just focus on the simplest reasoning, and it is still very important.
@@
One point to be emphasized here is that the color correlation with more nucleons must be weaker, and why such a hidden-color state associated with the entire nucleus appears. The possibility is that the more nucleons this state is related to, the bigger the number of such states will be. This requires further theoretical confirmation. In Ref. Bakker14 , it is shown that, the number of the hidden-color states with 9 quarks is much more than the one with 6 quarks Bakker14 . In a nucleus, the hidden-color states with various nucleon numbers may exist, but the most important may be the two-body correlation (the strongest), and the overall correlation caused by the hidden-color states with the nucleus number $A$ (the most possibilities). This is exactly what the nuclear experiments have been found out to be.
@@
Since there are many hidden-color states in the nucleus, a possible result will occur. These hidden-color states will have quantum interference with each other, and if destructive interference occurs, the long-range correlation may disappear and the magic number appears. If constructive interference occurs, a large deformation will occur. This is what to be focusing on in the future. If the magic numbers can be explained based on this result, it will further prove the correctness of this conclusion here.
@@
Just as Bakker and Ji said:“ Nuclear systems are identified as color-singlet composites of quark and gluon fields, beginning with the six-quark Fock component of the deuteron. An immediate consequence is that nuclear states are a mixture of several color representations which cannot be described solely in terms of the conventional nucleon, meson, and isobar degrees of freedom: there must also exist hidden-color multi-quark wave-function components, i.e., nuclear states which are not separable at large distances into the usual color-singlet nucleon clusters.” The nucleons or their combinations may have the color-neutron component (0,0), but they can not the (0,0) representation. In this Letter, I give the first evidence that this hidden-color state should exist due to the existence of the Cd puzzle and its explanation with the SU3-IBM.
@@
Second the long-range correlations are not (0,0). This long-range correlations lead to the deformations of the nucleus. If they are not (0,0), it means that it cannot be spherical. This is exactly what the Cd puzzle is trying to reveal. This conclusion is also supported when the Cd puzzle is verified by the SU3-IBM. In the new spherical-like $\gamma$-soft mode, the SU(3) third-order Casimir operator is as important as the second-order one. This is a new kind of deformation. In Ref. wang23 , the shape phase transition from the prolate shape to the oblate shape in the Hf-Hg region is studied, and the ratio of the SU(3) second-order and third-order Casimir operators is shown in Fig. 1. This value decreases approximately linearly as the number of nucleons increases (the number of bosons decreases). Recently it is also found that the boson number odd-even effect Fortunato11 ; Zhang12 of the 196-204Hg really exists, in which the SU(3) third-order interaction plays the main role Wang . This means that when the valence nucleons increase, the two-body residual interaction may not dominate. This is not quite the usual mean field viewpoint. I guess that the multiple possibilities of the color confinement in the nucleus in equation (1) let the low-energy two-body interactions interfere destructively, reducing it. The appearance of the magic numbers may arise from the complete quantum interference between the hidden-color states. For the Cd puzzle can be well explained by the SU3-IBM, the hidden-color state in equation (1) is much more important than expected.
@@
This hypothesis is strongly supported by the EMC effect, which needs to be further verified. The Cd puzzle is the direct consequence of this hypothesis. Although this explanation is easy, it fits with the results of the experiments. The conclusion that spherical nucleus does not exist can be confirmed by the Cd puzzle experiments and can be supported by other surrounding nuclei, such as Te, Pd nuclei Garrett18 . The results of these experiments can be well described by the SU3-IBM Wang22 ; WangPt ; Wang24 , which further supports this hypothesis.
@@
The result is somewhat unexpected. For the first time, it explicitly points out a clear correlation between the low and high energy behaviors of the nucleus, and while the true reason for the correlation is still unknown, it does exist. So this hypothesis is supported by both experiment and theory. The role of the quarks and gluons in the nucleus, though not clear, exists and works at low energy excitation. This has been overlooked in previous studies. Of course, this is inevitable. Because the change of the quarks and gluons in the nucleus is the behaviors found in the high-energy region, it is difficult to expect that the EMC effect can affect the low-energy nuclear structure. This requires support from the experiments. The Cd puzzle, as well as its SU3-IBM description, becomes the possible evidence.
@@
The experimental study of the short distance interactions between nuclei is the key. I expect to find some experimental phenomena that are sensitive to the nucleon number $A$ in the nucleus Sargsian14 and correspond to the sensitivity of the shape of the nucleus to the nucleon number $A$. For the short distance interactions, the effect induced by the hidden-color states may be only a part, so it may not be easy to distinguish experimentally.
@@
In conclusion, a simple explanation for the absence of the spherical nuclei is given based on the EMC effect, the short-range strong correlations, the Cd puzzle and other new experimental results, the calculation of the SU3-IBM and the hidden-color state. It can be obtained that the Cd puzzle is the result of the color confinement in the nucleus. If the hidden-color states exist, the spherical nucleus can not appear. It is clearly pointed out that the color degree of freedom is necessary even in low energy excitation of nuclear structure. The collectivity of the nucleus results from the hidden-color states. The correlations between the low-energy and high-energy behaviors in the nucleus does exist, and this is the first time to be found here, but what it is requires more further studies.
In this Letter a new theoretical framework to understand the nuclear forces, especially the short distance quark-gluon interactions, is established phenomenologically and a possible big picture for the whole nuclear physics is drawn. Since this explanation is very simple and rigorous, the new connection deserves serious consideration. It does not conflict with previous effective field theories. The possibility of the color degree of freedom between the nucleons is considered here. The SU3-IBM supports this possibility and gives a guide for future theoretical developments. In future, the possible results of this new perspective will be discussed.